wo tiny, spherical water drops, with identical charges of −3.20 ✕ 10−17 C, have a center-to-center separation of 1.10 cm.

(a) What is the magnitude of the electrostatic force acting between them?
___________________N

(b) How many excess electrons are on each drop, giving it its charge imbalance?
__________________ electrons
Answer #1

a)

magnitude of electrostatic force = k * q^2/d^2

magnitude of electrostatic force = 9 *10^9 * (3.2 *10^-19)^2/0.011^2

magnitude of electrostatic force = 7.62 *10^-24 N

the magnitude of electrostatic force is 7.62 *10^-24 N

b)

number of excess electrons = -3.2 *10^-17/(-1.6 *10^-19)

number of excess electrons = 200

the number of excess electrons is 200